### v1

# LLMs as verification oracles for Solidity

Abstract
Ensuring the correctness of smart contracts is critical, as even subtle flaws can lead to severe financial losses. While bug detection tools able to spot common vulnerability patterns can serve as a first line of defense, most real-world exploits and losses stem from errors in the contract business logic. Formal verification tools such as SolCMC and the Certora Prover address this challenge, but their impact remains limited by steep learning curves and restricted specification languages. Recent works have begun to explore the use of large language models (LLMs) for security-related tasks such as vulnerability detection and test generation. Yet, a fundamental question remains open: can LLMs serve as verification oracles, capable of reasoning about arbitrary contract-specific properties? In this paper, we provide the first systematic evaluation of GPT-5, a state-of-the-art reasoning LLM, in this role. We benchmark its performance on a large dataset of verification tasks, compare its outputs against those of established formal verification tools, and assess its practical effectiveness in real-world auditing scenarios. Our study combines quantitative metrics with qualitative analysis, and shows that recent reasoning-oriented LLMs can be surprisingly effective as verification oracles, suggesting a new frontier in the convergence of AI and formal methods for secure smart contract development and auditing.
1 Introduction
Since deployed smart contracts are immutable and often manage substantial economic assets, ensuring their correctness and security is paramount. Vulnerabilities and flaws in business logic have already resulted in losses exceeding \$6 billion [16], underscoring the urgent need for safer and verifiable code.
For Solidity, the most adopted smart contract language, there exist lots of bug detection tools [44, 49], which focus on finding common vulnerability patterns (e.g. re-entrancy, overflows, locked ether). However, it has been shown that the vast majority of the losses due to real-world attacks are caused by logic errors in the contract code [16]. This motivates the need for formal verification tools, able to verify (or falsify) custom properties concerning the business logic of a contract. Several verification tools have been developed for this purpose, most notably SolCMC [3], shipped with the Solidity compiler, and the Certora Prover [25], a leading tool in the auditing industry.
While, in general, formal verification tools can significantly enhance the security of smart contracts, they also present several limitations:
  • properties need to be written in a formal specification language, which
    1. (a)
      requires a considerable manual effort,
    2. (b)
      restricts the set of expressible properties;
  • for efficiency reasons, they often introduce unsound approximations of the smart contract semantics (e.g., they abstract from the gas mechanism, or neglect semantical corner cases such as coinbase and selfdestruct transactions);
  • despite such approximations, inherent issues of formal verification (e.g. state explosion, undecidability of non-linear arithmetic) are unavoidable.
These limitations are not merely theoretical, but also have a practical impact on current Solidity verification tools. Both SolCMC and the Certora Prover have been shown to be unsound (i.e., they may report a property as verified when it is not) [9], and relevant classes of properties (e.g., strategic ones) remain beyond their scope [7]. Such shortcomings highlight the need for alternative techniques that can complement symbolic formal verification.
Recently, there has been a huge rise in works that adopt Large Language Models (LLMs) [11] for a variety of tasks in smart contract analysis, including common vulnerabilities detection [17, 18, 30, 47], test generation [6], and specification generation [28, 31]. While LLMs have been shown to perform reasonably well on such tasks, it remains unexplored whether they can also verify custom, contract-specific properties. This challenge is considerably more complex, since the properties to verify can be arbitrary and are strictly tied to the specific contract under analysis. Therefore, mere pattern recognition is not sufficient; a certain degree of reasoning capabilities is required. Recent models such as GPT-5 (released on August 7th, 2025) have claimed to show human-like reasoning capabilities; however, whether these capabilities transfer to smart contract verification has not been explored yet.
If proven to be viable, the adoption of LLMs in the context of smart contracts verification would be quite appealing. Indeed, in contrast to formal verification tools, LLMs have the benefits of: (i) being able to work with arbitrary custom properties written in natural language, (ii) having a lower learning curve, (iii) adopting a knowledge-based and pattern-driven problem-solving approach that complements the symbolic reasoning approach used by formal methods [40]. On the other hand, however, LLMs present significant downsides compared to formal verification: well-known weaknesses include hallucinations, non-determinism, and lack of mathematical rigor.
In this work, we investigate the viability of using LLMs for smart contract verification. In particular, we consider the following research questions:
RQ1)
Given a contract and a custom property written in natural language, are LLMs able to correctly determine whether the property holds or not?
RQ2)
Analysis of the output:
  • a)
    When LLMs correctly guess the truth value of a property, are they also able to produce a correct explanation or counterexample that substantiates their answer?
  • b)
    When LLMs incorrectly guess the truth value of a property, what are the causes of failure?
RQ3)
What are the strengths and weaknesses of LLMs compared to mainstream Solidity verification tools?
To rigorously address these research questions, it is essential to operate in a controlled environment where the ground truth of the considered properties can be manually established. To further investigate whether LLMs can be useful in the wild, we consider a fourth research question:
RQ4)
Can LLMs be effectively used in real-world auditing scenarios?
This paper addresses the previous research questions by providing the following key contributions:
  • For RQ1-3, we constructed the largest public dataset for Solidity verification, to the best of our knowledge. The dataset defines 2034 verification tasks distributed across $5$ paradigmatic use cases, each one implemented in several mutations containing subtle bugs or alternative behaviours. The dataset contains a wide range of properties, encompassing function specs, state and transition invariants, and also complex strategic properties that are beyond the scope of existing symbolic verification tools.
  • For RQ1, we evaluated GPT-4 and GPT-5 over our dataset. Results show that GPT-5 consistently achieves over $85\%$ across all prediction metrics and use cases, outperforming GPT-4 by a margin of $25\%-30\%$.
  • For RQ2, we manually analysed the explanations and counterexamples produced by the LLMs to justify their classifications. This analysis confirmed that both models properly interpreted the natural-language properties. We then assessed the correctness (i.e., absence of logical errors) and completeness of the explanations, finding a clear correlation between GPT-5’s high prediction metrics and its stronger reasoning capabilities. In contrast, GPT-4 often produced vague or erroneous arguments leading to hallucinations.
  • For RQ3, we compared LLMs and the state-of-the-art symbolic verification tools SolCMC and Certora, finding that GPT-5 achieves higher prediction metrics than the two tools on their respective expressibility sub-datasets, also behaving consistently well on the set of the (often more complex) properties not expressible in the two tools.
  • For RQ4, we considered Safe and InfiniFi, two projects recently audited by Certora [15]. We ran GPT-5 over the properties addressed in the two reports, showcasing the ability of GPT-5 to correctly classify most of them. We further demonstrated how GPT-5 enhances verification auditing by detecting inconsistencies between property descriptions and formal specifications, and substantiating violations by producing detailed counterexample traces.
Overall, our analysis shows that recent models such as GPT-5 can be effectively used to support smart contract auditing. To foster reproducibility, we release our dataset, the scripts needed to run the experiments, and the complete experimental results in a public (anonymised) repository [2].
2 Background
This work lies at the intersection of formal verification and AI–based code analysis. We provide below the an essential background on these topics.
2.1 Symbolic verification tools for Solidity
Several verification tools for Solidity have been proposed in recent years (see Section 9). Among them, two have achieved industrial adoption: SolCMC [4], a symbolic model checker integrated with the Solidity compiler, and the Certora Prover [25, 10], one of the leading tools in security auditing of smart contracts.
In SolCMC, properties are expressed through assertions embedded in Solidity code. These assertions specify state invariants that must hold in all reachable contract states. SolCMC translates the instrumented contract into a set of logical constraints, which are then fed to a SMT/CHC solver (Z3 [34] or Eldarica [23]).
Unlike SolCMC, Certora decouples the specification of the properties from the contract code, providing a domain-specific language (CVL) for specs. A CVL spec defines constraints on the execution of a contract, and assertions that must be true in all states satisfying the given constraints. Certora compiles the Solidity contract and the CVL spec into a logical formula, and relies on SMT solvers to determine if the spec is satisfied in all contract states.
Both tools address the state explosion problem inherent in symbolic verification by approximating the Solidity semantics (e.g., both of them abstract from gas costs). These approximations, however, may yield false negatives (e.g., a property holds but the tool classifies it as violated), or false positives (the property is violated, but the tool reports it as satisfied). Assessing the reliability of a classification then requires a deep expertise on both the Solidity semantics and the specific approximations made by the tool, which in practice leads to a costly, iterative process of refining the specifications.
2.2 Large Language Models for reasoning
Large Language Models are trained to predict the next token in a sequence, effectively capturing statistical regularities from massive text corpora. At their core, these models do not reason in the human sense: they generate continuations based on learned patterns rather than explicit symbolic inference. Nevertheless, suitable prompting strategies can produce behaviors that resemble logical or step-wise reasoning, enabling models to address multi-step problems beyond surface-level pattern matching. In this context, reasoning refers to the model’s ability to decompose a task into intermediate steps, evaluate partial conclusions, and synthesize them into a final answer. A consistent finding is that prompting models to think in steps improves accuracy on multi-step math, logic, and symbolic tasks. In particular, Chain-of-Thought (CoT) prompting, eliciting brief intermediate rationales, boosts performance across arithmetic, commonsense, and symbolic reasoning benchmarks, especially for larger models [46].
A closely related variant is Zero-shot CoT, which adds a simple instruction to reason step-by-step without providing exemplars. This single instruction has been shown to markedly improve zero-shot results on diverse reasoning benchmarks [27]. Prior work shows that promising results can be achieved by prompting the model with a Zero-shot CoT with hidden rationales (i.e., the reasoning is performed but not revealed verbatim): this step-wise prompting can improve reliability without requiring few-shot demonstrations [27].
3 Dataset design
Our research questions require both a systematic analysis (RQ1–3) and coverage of real-world scenarios (RQ4). Following common practice [17, 21, 47], we therefore combine complementary methodologies. For RQ1–3, we construct a controlled dataset of paradigmatic use cases and verification tasks, for which the ground truth can be manually established. We describe in this section the design of the dataset, in the following section the methodology used to perform the experiments for RQ1-3, and in Section 8 the related discussion for RQ4.
Table 1: Dataset metrics (lines of code averaged over number of versions).
Use case LoC #Mut #Prop       Solidity features
Bank 28 17 27 ETH transfer, mappings, linear arith.
Vault 55 11 38 ETH transfer, time, state machine
PriceBet 73 16 42 ETH transfer, external oracle
PaymentSplitter 145 10 17 ETH transfer, loops, non-linear arith.
LendingProtocol 303 9 35 ERC20 tokens, mappings, non-linear arith.
Our dataset comprises 5 use cases, covering different Solidity features and different levels of complexity (Table 1):
  1. 1.
    Bank: this is a minimal wallet contract that allows users to deposit and withdraw ETH. Its simplicity allows us to design properties whose truth may depend on subtle aspects of the Solidity semantics, e.g. contract-to-contract calls, gas, reentrancy, and selfdestruct transactions. The properties in the dataset include function specs (e.g., a non-reverting call to a certain function produces a certain effect on caller and callee), state invariants (e.g., the sum of all credits does not exceed the contract balance), metamorphic properties (e.g., two different sequences of transactions produce the same effect on the contract state), and strategic properties (e.g., there exists a sequence of transactions that produce some desired effect on the contract state — such as no ETH remains frozen in the contract).
  2. 2.
    Vault: this is a contract allowing the owner to deposit and withdraw ETH, with a mechanism to cancel pending withdrawal requests through a recovery key. Its implementations require time constraints and an automaton-like behaviour (i.e., certain actions can only be performed in certain states). The properties now also predicate about the immutability of contract variables (e.g., a certain address is the same in every contract state — or even within the execution of a function), and liveness properties with fairness assumptions (e.g., “if $A$ knows both the owner and recovery key, and no one else knows the recovery key, then in every fair trace $A$ is able to eventually transfer any fraction of the contract balance, while no $B\neq A$ can do the same”).
  3. 3.
    PriceBet: this is a two-players bet on the price of a token, which is queried from an external oracle [5]. The properties in the dataset explore additional aspects such as assumptions about external contracts and MEV attacks (e.g., an adversary can frontrun another player’s transaction to win the bet).
  4. 4.
    PaymentSplitter: this is an OpenZeppelin template that allows a set of shareholders to pull payments from a contract proportionally to their shares [37]. It allows us to explore the impact of non-linear arithmetic (used to determine the releasable ETH for each payee) and loops (used in the constructor to initialize the shareholders data).
  5. 5.
    LendingProtocol: this is the most complex contract in our controlled dataset, as it implements a subset of the functionality of the Aave protocol [1] (with some simplifications in order to keep the ground truth reliable). Its allow us to further explore non-linear arithmetic, integer truncations, and economic properties (e.g., certain actions produce certain gains for the involved users).
The mutations of each use case are constructed manually, by introducing logic errors or divergent behaviours that may affect the validity of some the given properties. We provide each use case with a set of properties, written in natural language, which are common to all the mutations. Some properties are given in slightly different versions, to assess how verification is affected by different assumptions on the context (e.g., sometimes we assume that an address has a receive function that just accepts all ETH, in order to verify if ETH transfers work as intended at least in this simple case).
For each verification task — i.e., pair (property, mutation) — we provide the ground truth, i.e. a boolean telling whether the property holds or not for that mutation. The first three use cases (Bank, Vault and PriceBet) are simple enough for us to enable us to construct this ground truth manually (despite the apparent simplicity, however, there are several cases for which a certain degree of expertise is required, e.g. when the ground truth is affected by quirky semantical aspects such as reentrancy, selfdestruct transactions, or gas passed along with low-level calls). For the most complex use cases (PaymentSplitter and LendingProtocol), besides manual analysis, we have consolidated the ground truth by providing Hardhat PoCs for the verification tasks where the property is violated (the code of these PoCs is in the dataset). The verification tasks for which we were not sure about the ground truth are not included in the experiments.
There are two key remarks in the design of properties. First, we carefully crafted the natural language specifications to be as precise as possible, thereby reducing the risk of LLMs giving wrong answers because of misunderstandings. The qualitative analysis of the LLMs outputs (in particular, the coherency metrics, see Section 6) shows that the LLMs understood our specifications with high confidence. Second, we ensured a balanced range of complexity across properties, avoiding biases introduced by verification tasks that are either too easy or too difficult. Moreover, our dataset includes both properties expressible by symbolic verification tools (e.g., function specs, state and transition invariants) and properties beyond their expressive power, allowing us to assess the effectiveness of LLMs on tasks that remain out of reach for existing symbolic verification tools.
Symbolic verification tools.
To compare LLMs against symbolic verification tools, we focussed on SolCMC and Certora, for the reasons outlined in Section 2. For this purpose, we translated the natural-language properties into formal specifications for both tools and included them in the dataset. Such encodings were not always feasible: SolCMC is limited to state invariants, while CVL cannot express certain classes of properties, such as strategic ones. Hence, comparison with GPT-5 were performed on the relevant subsets of the dataset (see Section 7).
A crucial point here is that writing these formal specifications was not a one-shot process, but an iterative one. In particular, even when a property was true and its initial specification was a faithful encoding of the natural-language statement, the tool often failed to verify it. This issue arose more frequently with Certora, for two main reasons. First, Certora uses conservative over-approximations of contract behaviour, sometimes considering cases that are impossible in practice. For example, it may assume that a contract can call itself as a sender, even when such behavior is excluded by the code. In these situations, we had to introduce additional technical assumptions to exclude spurious behaviors that falsify the property. Second, some true properties require auxiliary invariants that the tool cannot automatically infer. In these cases, we supplied the missing invariants as part of the specification, effectively serving as proof hints to guide the verifier toward establishing the property.
4 Experimental setup for LLMs evaluation
The primary focus of this study is GPT-5, the most recent LLM released by OpenAI, which incorporates a dedicated reasoning module (GPT-5 Thinking) designed to address complex problems. As a baseline for comparison, we use GPT-4, which lacks such a reasoning component. To evaluate both models, we developed a Python script that takes as input a set of verification tasks, queries the LLMs via the OpenAI APIs, and collects their outputs for analysis.
Each query is constructed using a carefully crafted prompt that follows a zero-shot, chain-of-thought style. The prompt (see Appendix 0.A) requires the model to: 1 determine whether a given property holds across all reachable contract states; 2 explain the reasoning process leading to this decision; and 3 if the property does not hold, provide a counterexample in the form of a transaction sequence that violates it. We additionally allow the LLMs to respond with “unknown” when uncertain, although in practice we observed that both GPT-4 and GPT-5 rarely exercised this option. To ensure a fair comparison between LLMs and formal verification tools in RQ3, we made explicit in the prompt certain common abstractions: in particular, we instructed GPT to neglect gas costs: e.g., when an address $A$ sends a transaction to withdraw 1 ETH from the contract, we want GPT to assume that $A$’s balance increases by exactly 1 ETH, without accounting for the gas spent be $A$ to execute the transaction. This abstraction is consistent with the behaviour observed both in SolCMC and Certora.
To mitigate bias toward simpler tasks, we did not feed the models with the entire dataset. Rather, we sampled verification tasks from the three simpler use cases, while retaining the complete set of tasks from the others. The resulting sub-dataset (hereafter, referred to as $\mathrm{DS}$) contains 667 verification tasks, and it is nearly balanced, featuring $\sim\mbox{1:1}$ ratio between simpler and complex use cases, and a comparable $\sim\mbox{1:1}$ ratio between positive and negative tasks.
5 RQ1: LLMs metrics on the controlled dataset
Table 2: Quantitative analysis for GPT-4 (in gray) and GPT-5 (in cyan).
Use case Tot. UNK TP TN FP FN Accuracy Precision Recall Specificity F1 Score
Bank 121 0 39 47 21 14 71% 65% 74% 69% 69%
Vault 106 4 37 30 25 10 66% 60% 79% 55% 68%
PriceBet 158 4 39 55 34 26 61% 53% 60% 62% 57%
PaymentSplitter 160 2 43 44 41 30 55% 51% 59% 52% 55%
LendingProtocol 122 1 52 28 11 30 66% 83% 63% 72% 72%
OVERALL (GPT-4) 667 11 210 204 132 110 63% 61% 66% 61% 63%
Bank 121 0 50 65 3 3 95% 94% 94% 96% 94%
Vault 106 0 45 53 3 5 92% 94% 90% 95% 92%
PriceBet 158 0 60 84 8 6 91% 88% 91% 91% 90%
PaymentSplitter 160 0 68 70 15 7 86% 82% 91% 82% 86%
LendingProtocol 122 0 79 37 3 3 95% 96% 96% 93% 96%
OVERALL (GPT-5) 667 0 302 309 32 24 92% 91% 93% 91% 92%
Table 2 reports the experimental results of GPT-5 and GPT-4 on dataset $\mathrm{DS}$. GPT-5 achieves an F1 score above $86\%$ in all use cases ($92\%$ overall), outperforming GPT-4 on all of them with a performance gap of $25\%-30\%$ on every prediction metric. This result aligns with prior findings on GPT-4’s limited logical reasoning abilities [29]. The most challenging use case was PaymentSplitter, where GPT-4 performs only slightly better than random guessing, while GPT-5 still maintains performances above 80% across all metrics. We observe that GPT-5 has a very low rate of FP, which is a desirable soundness property of verification tools. We elaborate in Section 6 the main causes of FPs and FNs.
We were quite surprised by GPT-5’s performance on LendingProtocol, the most complex use case in our dataset. We elaborated some explanations for this result (see also Section 6). First, to address its complexity, we were more conservative than in the other use cases when crafting the mutations, to minimize the risk of errors in the ground truth. So, here the mutations do not exhibit the diversity of behaviours that we reached in the other use cases: this results in several properties sharing the same truth for most mutations. For the same reason, several properties we devised for LendingProtocol are function specs for the various operations (which are easire to check for the ground truth), while there are less strategic properties, which are more difficult to assess manually, and arguably also for the LLMs. Additionally, this use case deals with tokens instead of ETH: to avoid dealing with malicious token implementations, in many properties we added the assumption that they are standard ERC20 tokens: this ruled out reentrancy, which was one of the causes of failure of GPT-5 in other use cases. We also observed that GPT-5 is very good at reasoning about non-linear integer arithmetic, which was required to solve many of the verification tasks.
6 RQ2: Analysis of LLMs explanations
Table 3: Qualitative analysis. In gray, results for GPT-4 on the Bank use case. In cyan, results for GPT-5 on all use cases.
TP + TN FP + FN
Use case # Correct Complete Coherent # Correct Complete Coherent
Bank 86 78 85 84 35 2 13 32
Bank 115 115 105 114 6 1 4 5
Vault 54 54 49 53 8 5 7 4
PriceBet 74 73 66 74 10 7 5 8
PaymentSplitter 13 12 11 13 4 0 0 4
LendingProtocol 17 17 16 17 3 2 2 1
We now study the quality of the explanations and of the counterexamples provided by the LLMs. We manually analyzed answers of GPT-4 and GPT-5 (to, respectively, 121 and 304 tasks), evaluating them under three criteria: coherence, completeness, and correctness. By coherence, we mean that the property has been properly understood by the LLM: we assess this by checking that the conclusions indeed imply that the property is verified/falsified, and that, if some hypotheses have been used, they match the ones in the property. By correctness, we mean that the explanation does not contain logical errors: for True-answers, each implication “premise $\implies$ conclusion” must be indeed logically valid; for False-answers, we additionally check that each step of the counterexample is executable in the given states. By completeness, we mean, for True-answers, that the explanation uses all the assumption mentioned in the property that are necessary for it to hold; for False-answers, that the counterexample contains all the necessary steps. Overall, correct and complete counterexamples can be translated into executable PoCs.
For each answer, we mark each criterion with 0 or $1$, and add annotations whenever necessary. Table 3 reports the results. Although this analysis necessarily present a certain degree of subjectivity, it helps better understanding how faithfully the metrics for RQ1 represent the reasoning capabilities of LLMs.
To compare GPT-4 vs. GPT-5, we focussed on the Bank contract, since its simplicity allowed us to exhaustively evaluate all the answers. As we can see, coherence is rarely an issue, suggesting that both models are rather able to correctly interpret the property; this is further supported by the fact that slight variations of the properties (e.g., different assumptions) often lead to the expected explanations and counterexamples. The main cause of false results for GPT-4 is incorrectness. GPT-5, on the contrary, performs reasonably well on all criteria. In case of wrong answers, the cause distributes rather equally among the three criteria; in case of correct answers, the main weak point is completeness.
In order to answer to RQ2b, we analyzed the causes of incorrect answers. By manually inspecting the verification tasks where FPs are produced, we noted that they have a few common causes: 1 low-level calls triggering ETH transfers; 2 selfdestructtransactions increasing the contract balance; 3 integer overflows (mostly, in PaymentSplitter). We note, however, that GPT-5 is not consistently wrong when dealing with these features: rather, it answers correctly most of the times. A typical cause for FNs is the use of transfer to send ETH: here, GPT-5 often misses that this operation does not carry enough ETH to allow the callee to perform a selfdestruct.
7 RQ3: LLMs vs. formal verification tools
Table 4: Expressibility (left) and Coverage (right, in parentheses) percentages.
Use case # SolCMC Certora GPT-4 GPT-5
Bank 121 62.81   (61.98) 89.26   (89.26) 100   (100) 100   (100)
Vault 106 26.42   (26.42) 54.72   (54.72) 100   (96.23) 100   (100)
PriceBet 158 47.47   (46.84) 56.33   (56.33) 100   (97.47) 100   (100)
PaymentSplitter 160 43.75   (18.13) 87.50   (87.50) 100   (98.75) 100   (100)
LendingProtocol 122 4.10   (0) 81.15   (77.05) 100   (99.18) 100   (100)
OVERALL 667 38.08   (30.89) 74.06   (73.31) 100   (98.35) 100   (100)
As discussed in Section 3, only certain properties can be expressed in the specification languages of SolCMC and Certora. We define expressibility as the percentage of verification tasks that can be expressed in the given tool (for LLMs, this is, by default, $100\%$). We also measure the coverage, i.e. the percentage of tasks for which the tool returns True or False (instead of Unknown or timeouts).
[…]
To ensure a fair comparison of the LLMs metrics in Section 5, which cover the entire dataset $\mathrm{DS}$, with those of SolCMC and Certora, we define two sub-datasets corresponding to the expressible verification tasks of the two tools: we refer to them as $\mathrm{\mathrm{DS}_{|SolCMC}}$ and $\mathrm{\mathrm{DS}_{|CVL}}$, respectively. To study how LLMs perform on the tasks not expressible by SolCMC and Certora, we define the complementary sub-datasets $\mathrm{\mathrm{DS}_{|\neg SolCMC}}:=\mathrm{DS}\setminus\mathrm{\mathrm{DS}_{|SolCMC}}$} and $\mathrm{\mathrm{DS}_{|\neg CVL}}:=\mathrm{DS}\setminus\mathrm{\mathrm{DS}_{|CVL}}$}.
Table 5: Results by sub-datasets (all use cases).
Dataset Tool Tot. UNK TP TN FP FN Accuracy Precision Recall Specificity F1 score
$\mathrm{\mathrm{DS}_{|CVL}}$ Certora 494 5 196 179 58 56 77% 77% 78% 76% 77%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-5 494 0 238 214 24 18 91% 91% 93% 90% 92%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-5 173 0 63 96 9 5 92% 88% 93% 91% 90%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-4 494 6 173 149 86 80 66% 67% 68% 63% 68%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-4 173 5 37 55 46 30 55% 45% 55% 54% 49%
$\mathrm{\mathrm{DS}_{|SolCMC}}$ SolCMC 254 48 72 95 14 25 81% 84% 74% 87% 79%
$\mathrm{\mathrm{DS}_{|SolCMC}}$ GPT-5 254 0 110 119 19 6 90% 85% 95% 86% 90%
$\mathrm{\mathrm{DS}_{|\neg SolCMC}}$ GPT-5 413 0 191 191 14 17 92% 93% 92% 93% 92%
$\mathrm{\mathrm{DS}_{|SolCMC}}$ GPT-4 254 1 84 84 53 32 66% 61% 72% 61% 66%
$\mathrm{\mathrm{DS}_{|\neg SolCMC}}$ GPT-4 413 10 126 120 79 78 61% 61% 62% 60% 62%
Table 5 shows the results of the analysis. Under all metrics, GPT-4 performs considerably worse than SolCMC and Certora on their respective datasets, while GPT-5 performs considerably better than Certora on all metrics, and better than SolCMC on all metrics except specificity (and it wins only by a narrow margin on precision). This is due to the low number of FPs in SolCMC. Remarkably, GPT-5’s performance is consistent across the restricted sub-datasets and their complements. This indicates — perhaps surprisingly — that the properties not expressible in the two verification tools are not inherently harder for GPT-5.
Besides correctly identifying the validity of a property, the value of a verification tool lies in the guarantees it can provide about the answer (i.e., a counterexample or proof certificate). In Section 6, we analyzed the answers of GPT-4 and GPT-5. Here, we discuss how they compare with those of SolCMC and Certora. Both tools, when returning True, do not provide any explanation: the guarantees just lie on their correct implementation and on their symbolic interpretation of the Solidity semantics (but, as observed before, some approximations they make are unsound). When returning False, SolCMC tries to return a sequence of transactions that brings the contract in a state where the invariant is violated. In most experiments, however, it did not produce any counterexample; moreover, it may return FNs in certain cases in which it abstracts from the Solidity semantics [42]. Certora, on the contrary, over-approximates the set of reachable states: its counterexamples are states which violate the property, but they are not guaranteed to be reachable, thereby producing spurious counterexamples [13]. This explains the high number of FNs in Table 5.
8 RQ4: LLMs for verification auditing in the wild
In order to assess whether GPT-5 can be useful in real-world auditing scenarios, we selected the two most recent formal verification auditing reports made available by Certora at the time of writing: Safe v1.5.0 and infiniFi Protocol [15]. The Safe contract is a smart account wallet [20], while infiniFi is a protocol that offers yield-bearing stablecoin deposits [24].
The reports contain properties written in natural language, each of which is associated to a formal specification written in CVL. We feed GPT-5 the properties written in natural language, and ask it to verify them with the same prompt used in the other experiments (i.e., the one in Appendix 0.A). We stress that, since GPT-5 knowledge cut-off is September 30th 2024 [36], we can exclude that the model uses information found in the reports. This is further validated by the fact that GPT-5 often disagrees with Certora.
In general, we cannot determine the ground truth of the properties, as these contracts consist of thousands of lines of code, besides those in external dependencies. Hence, reliably establishing the validity of their properties is rather unfeasible, forcing us to recur to a more empirical approach.
We carefully analysed GPT-5 answers, with particular attention to properties over which GPT-5 disagrees with Certora. First, we examined the CVL specification, to verify whether it is consistent with the property written in natural language; if not, we rephrased the property accordingly, and re-ran GPT-5 over the newly formulated property. If the cause of disagreement is not an inconsistency between the formal specification and the property description, we then analysed the counterexample provided by one of the two tools. In particular, since the states that Certora provides as counterexamples lie on an over-approximation of the set of reachable states, we had to ensure that the returned state is indeed reachable, in order to confirm the violation. In order to do that, we paired manual inspection with new additional queries to GPT-5, feeding it the Certora’s state and asking it to provide a trace that starts from the deployment of the contract and leads to the state under analysis.
We summarise below the results of our experiments (available in [2]), and then discuss our findings. An extensive report of the results is in Appendix 0.B.
Summary.
The Safe report contains 25 properties, all with reported True status according to Certora, while the Infinifi report contains 47 properties, 8 of which are reported to have False status and 41 True status, according to Certora.
We summarize our observations as follows:
  • Property/Spec mismatch: In 12 cases, GPT-5 and Certora disagree due to a mismatch between the property description and the CVL spec, either due to assumptions made in the spec but missing in the property, or due to typos (e.g. $\leq$” in the spec but “<<” in the property). In all these cases, manual inspection confirms that GPT-5 results are correct w.r.t. the property description. If we rectify the properties to match the CVL spec, GPT-5 returns results consistent with Certora’s.
  • Munged contract: In 6 cases, GPT-5 and Certora disagree due to the fact that Certora was run on a munged version of the contract, in which delegate calls were disabled. This was not specified in the report, and we had to inspect the CVL run to notice it. Manual inspection confirms that the violations found by GPT-5 indeed hold on the original contract. If we instruct GPT-5 to disregard delegate calls, it returns results consistent with Certora’s.
  • Cryptographic assumptions: In 1 case, GPT-5 returns Unknown as the validity of the property depends on a cryptographic assumption, which the Certora Prover makes internally [14]. If we add this assumption to the property, GPT-5 returns a result consistent with Certora’s.
  • Math errors: In 4 cases, GPT-5 makes errors involving numbers in scientific notation. Instructing it to always convert to decimal notation before performing computations was sufficient to prevent it from repeating the errors.
  • Missing violations: In 3 cases, GPT-5 does not spot a violation. In 1 case, it does not detect unstable rounding behaviors; in 2 cases, it overlooks hooks calls that can cause the violation (we note that, if we instruct GPT-5 to “Pay attention to hooks”, it can spot the violations).
  • Validating Certora’s cex: For each property claimed as False by Certora, we fed the returned state to GPT-5 and asked it to verify whether the state is reachable or not, and, if yes, to provide a trace that starts from deployment. In all cases except one, GPT-5 confirms that the state returned by Certora is indeed reachable, validating the violations found. In the remaining case, even manual inspection was not able to verify whether the state is indeed reachable; in this case, the correctness of the violation found by Certora is questionable. We refer to Section 0.B.2 for a detailed discussion.
Discussion.
We first note that, in most cases, GPT-5 returned answers that match what we believe to be the respective ground truths. In particular, in most of the cases in which a property is violated, GPT-5 was able to detect the violation and return a valid counterexample. Furthermore, we observed a powerful synergy between GPT-5 and Certora for the validation of violation counterexamples. Indeed, Certora returns states that are not guaranteed to be reachable (and the tool itself is not able to prove reachability, in general). GPT-5, on the other hand, is quite good at producing detailed traces that start from contract deployment to reach a given state.
Another positive role GPT-5 excels at is spotting inconsistencies between the description of properties in natural language and the formal specifications. This is a rather important aspect of the auditing process, as the quality of a report also depends on the correct explanations of what has been verified, precisely (readers are not supposed to investigate formal specifications themselves, but rather trust the property descriptions available in the report).
A weak point of GPT-5 is that it struggles with scientific notation, which causes it to fail to find vulnerabilities related to subtle mathematical dynamics.
We also observed that GPT-5 sometimes improves its answers when given targeted hints, such as paying particular attention to some features (e.g. hooks) or to some methods. This is not very different from how auditors use Certora, which often needs to be instructed to only analyse certain parts of the code, or to adopt simplifying assumptions without to avoid timeout, etc. This further validates that human guidance still plays a relevant role in the auditing process, regardless of the specific tool used.
9 Related work
Research on smart contracts analysis has traditionally followed two main directions: vulnerability detection and formal verification. The former focusses on finding bugs in contract code by leveraging recurring vulnerability patterns. In this setting, there exist large datasets of contracts tagged with their vulnerabilities that are used as benchmarks to evaluate new analysis techniques [38, 41, 45]. Formal verification instead focusses on proving that a contract satisfies specific properties of its business logic (or showing counterexamples otherwise). This approach is pursued through different perspectives, often related to the class of expressible properties. For example, VeriSolid [35] focusses on CTL properties; SmartPulse [39] on past-LTL properties; Solvent [8] on strategic properties of the form “there exists some transaction leading to a certain state”. Compared to SolCMC and Certora, however, these tools are not yet directly usable in the wild, as they typically support restricted fragments of Solidity. Unlike the vulnerability detection setting, no large datasets exist for benchmarking formal verification tools. The largest available dataset so far contains only 323 verification tasks [9]; whereas the dataset developed for this work comprises 2034 tasks.
LLMs offer a complementary path to contract analysis: thanks to their understanding of natural language and reasoning capabilities, they may help overcoming some of the limitations of formal verification tools. David et al. [17] explore the use of LLMs for smart contract auditing, by querying GPT-4 and Claude to detect the presence of a pre-defined vulnerability patterns (e.g. reentrancy, delegatecall injection, etc.). Using a benchmark of 52 compromised DeFi contracts, they show that GPT-4 and Claude accurately identify vulnerabilities in ${\sim}40\%$ of cases, albeit the high FP rate still requires human oversight. At the light of our results, we believe that GPT-5 would substantially improve the metrics of [17]. SmartGuard [18] addresses vulnerability detection by combining the retrieval of semantically similar code with chain-of-thought generation and a self-check architecture, validating their technique on the SolidiFI benchmark [19]. The work [30] fine-tunes ChatGPT on 7 logical vulnerabilities (e.g., price oracle manipulation, privilege escalation), and experiments on a dataset of audit reports. LLM-SmartAudit [47] employs a multi-agent conversational framework with a buffer-of-thought mechanism, addressing 10 types of common vulnerabilities (reentrancy, overflow, etc.). While the previous three works achieve high prediction metrics on their respective datasets, the use of different datasets for validation makes direct comparison challenging.
Compared to the previous works, ours takes a different perspective: rather than focussing on detecting known vulnerability types, we investigate whether LLMs are capable of reasoning about arbitrary, contract-specific properties, and whether they can provide convincing explanations for their answers.
During the construction of our dataset, we observed that crafting properties that precisely capture the desirable contract behaviors is quite challenging. To address this issue, some works have proposed LLMs as an aid to automate the generation of such properties. PropertyGPT [31] uses LLM-based in-context learning to generate properties from contract code, targeting in a custom specification language similar to Certora’s CVL. The work [28] also uses LLMs, targeting post-conditions in the language of solc-verify [22]. A limitation of both works is that they cannot deal with arbitrary properties, since after generation they still resort to symbolic tools for verification. As noted before, this rules out relevant classes of properties (e.g., strategic ones), which could provide more precise insights on the desirable contract dynamics. Indeed, it has been observed that “global” properties (such as metamorphic or strategic properties) usually yield a higher return of investment than “local” ones, such as function specs [48].
While our work only hints at the potential of combining LLMs with formal methods, other studies — in domains outside of smart contracts — have already shown that such integration can be quite effective in practice. For example, ESBMC-AI [43] integrates LLMs with bounded model checking to detect and repair common flaws. The work [26] show that LLMs can aid in formal verification tasks, such as inferring loop invariants and proving safety properties.
10 Conclusions
In this paper, we have investigated the use of LLMs for smart contract verification. Our systematic evaluation over a large dataset of verification tasks, as well as experiments on two real-world smart contracts audits, substantiate that state-of-the-art reasoning-oriented LLMs such as GPT-5 can be surprisingly effective at the task. Due to their immediacy and ease of use, this can significantly push forward the accessibility of verification to a wider audience, enhancing smart contracts security. We also observed great complementarity with formal verification tools, suggesting hybrid approaches as the next frontier.
While this work paved the way to the use of LLMs for smart contract verification, several research questions remain open. A key challenge is how to improve the certifiability of LLMs answers, thereby reducing FPs and FNs.
For False-answers, our current approach is to ask the LLM to generate natural-language counterexamples that illustrate the violation. While this provides auditors with useful intuitions about why a property is violated, it does not guarantee that the example is reproducible on the actual blockchain. To reduce such false negatives, we could instruct the LLM to provide a concrete PoC, e.g. in the form of an executable Hardhat script (which we leveraged for the manually-crafted PoCs in our dataset).
For True-answers, ideally we would like the LLM to produce a machine-checked proof that the property holds in some proof assistant (e.g., in Rocq, Lean, Isabelle). While we see this as a challenging long-term goal, there are already a few works that may help in reducing the gap towards it. The work [33] formalises an executable semantics of a fragment of Solidity in the Isabelle/HOL proof assistant. Leveraging such semantics, it is possible to craft machine-checked proofs for True-properties. While this is done manually in [33], the long-term goal would be to exploit the reasoning capabilities of recent LLMs to assist humans in writing such proofs — if not even generate them autonomously (in other domains, there is already ongoing research [32, 12]).
Acknowledgments
Work partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union – NextGenerationEU, and by PRIN 2022 PNRR project DeLiCE (F53D23009130001). The authors thank Francesco Stori and Niccolò Viale for contributing to the construction of the dataset.
Appendix 0.A Prompt
The text below is the Python string that defines the prompt within the script used to call the GPT API. The prompt file, including the script and other resources used for the experiments, as well as the resulting outputs, is publicly available in the project repository [2].
You are an expert in Solidity smart contracts and formal verification.
You will be given:
1. A Solidity smart contract.
2. A property to verify on the contract, expressed in natural language.
Your task:
- Analyze the contract carefully.
- Determine whether the contract ALWAYS satisfies the given property.
- If the property is FALSE, you MUST provide a concrete counterexample,
for instance showing a state and/or a sequence of transactions that
violate the property.
- If you are NOT SURE, respond ""UNKNOWN"".
Think step by step internally about whether the property holds,
but DO NOT include your reasoning in the output.
In your reasoning, consider the following technical assumptions:
- ignore gas costs. For example, if a user sends a transaction
that is reverted, ignore the gas paid by the user;.
- when in a property we refer to a ""user"" of the contract, we implicitly
assume that the user is not the contract under analysis itself;
- there is not (and there will never be) enough ETH in circulation to
cause an overflow to a uint variable storing the balance of an address.
Always respond ONLY in the following format:
ANSWER: [TRUE | FALSE | UNKNOWN]
EXPLANATION: <brief explanation citing relevant lines or functions>
COUNTEREXAMPLE: <if ANSWER=FALSE, provide one; otherwise write ""N/A"">
Smart Contract:
{code}
Property:
{property_desc}
Appendix 0.B RQ4: LLMs for verification auditing in the wild
We detail the results of our experiments, respectively for the Safe contract (Section 0.B.1) and the InfiniFi protocol (Section 0.B.2). The complete report of all the experiments performed can be found in the public repository [2].
0.B.1 Safe
The report contains 25 properties, all with reported True status according to Certora.
We carefully analyzed the properties for which GPT-5 returned an answer different from that returned by Certora. Such set contains 14 properties overall: 13 reported as False and 1 reported as Unknown. We investigated the causes of disagreement. We organize our findings as follows:
  • Property/Spec mismatch: In 7 cases, the property written in natural language did not match the formal specification written in CVL.
    • in 4 cases, the CVL specification assumes that the mentioned call does not revert; without such assumption, the properties would indeed be violated.
    • in 1 case, the CVL specification assumes the value of a given parameter; without such assumption, the property would indeed be violated.
    • in 1 case (“setSafeMethodChanges”), the property description is ambiguous: it mentions that the call to a method changes the value of a variable; however, what is guaranteed is only that the call to the method sets the variable to the new specified value, which can be equal to the previous value. The report contains another property which precisely expresses this condition (“setSafeMethodChanges”). The CVL specifications of the two properties are indeed exactly the same.
    • in 1 case (“handlerCallableIfSet”), the CVL specification assumes that a given handle is a dummy method. Furthermore, we need to make the assumption that the choosen handler selector does not collide with that of any other existing function.
  • Munged contract: GPT-5 has found 6 properties which are violated due to the possibility of executing delegate calls. We inspected the counter-examples found, and they are indeed correct. The reason why the Certora Prover states that these properties are satisfied is because the prover was actually run on a munged version of the contract, in which delegate calls are not allowed. This assumption was nowhere specified in the report.
  • Cryptographic assumptions: In 1 case, GPT-5 returned Unknown, since the property under analysis only holds if we assume that keccak256 values cannot collide; the Certora Prover internally makes this assumption [14].
For all these properties, we have introduced new corresponding properties, adding the assumptions previously not specified. On all these new properties, GPT-5 returned True, matching Certora’s outputs. Note that, in all these cases, we never changed the code of the contract, but only the property description.
0.B.2 InfiniFi
The report contains 47 properties, 8 of which are reported to have False status and 41 True status, according to Certora.
We analyzed the set of properties over which GPT-5 returned an answer different from that returned by Certora (which contains 9 properties), and the set of properties over which both Certora and GPT-5 returned False (which contains 3 properties)
We organize our findings as follows:
  • Certora True / GPT-5 False: In 5 cases, Certora reported True while GPT-5 reported False:
    • in 4 of these cases, there is a mismatch between the property description and the CVL specification. In particular:
        in 1 case the property description uses a strict inequality (>>), while the CVL spec uses a non-strict inequality ($\geq$)
        in 1 case the property description excludes reverting, while the CVL spec excludes both reverting and returning 0,
        in 2 cases the CVL spec excludes calls to methods that are not excluded in the property description
      In all these cases, the property as written in natural language is indeed violated, and the counterexamples returned by GPT-5 are correct. If we re-run GPT-5 after adding the respective assumptions to the property descriptions, it returns True on all cases.
    • in 1 case, GPT-5 commits a math error involving scientific notation (e.g. it states that $\tt{6e18/3e18=2e18}$}). We instruct GPT-5 how to deal with numbers in scientific notation adding to the prompt “Be careful: whenever you have to perform arithmetical computations concerning numbers in scientific notations (e.g. 1e10, 6.3e14, etc), you must FIRST convert them to decimal notation (e.g. 10000000000, 630000000000000), and only THEN perform arithmetical computations.” After this instruction, GPT-5 returns again False, however providing a correct counterexample. After inspection, we noted that the CVL specification makes an additional assumptions, i.e. that USDC prices are around 12 orders of magnitude greater than iUSDC prices. If we add this assupmtion to the property description, GPT-5 returns True.
  • Certora False / GPT-5 True: In 4 cases, Certora reports False while GPT-5 returns True:
    • in 3 cases, after we asked GPT-5 to produce a complete trace that leads to the state provided by Certora as a counterexample, GPT-5 returned a complete and correct trace, validating Certora’s counterexample. In 2 of these 3 cases, the trace involves hooks, which have probably been overlooked by GPT-5 on the first run. If we suggest GPT-5 to be careful about hooks (i.e. we add “Pay attention to hooks.” to the prompt), it returns False, providing a correct counterexample. In the third case, the property is violated due to unstable rounding implementation: GPT-5 limited math reasoning is probably the reason why it was not able to spot the violation.
    • in 1 case, after we asked GPT-5 to produce a complete trace that leads to the state provided by Certora as a counterexample, GPT-5 claimed that the state is not reachable. We manually inspected that state, and have indeed some reservations regarding the reachability of the state. The following explanations has necessarily to enter into technical details. The CVL spec, in the precondition checks the total reward not through point.totalRewardWeight, but through totalRewardWeight(), which is equal to point.totalRewardWeight.mulWadDown(slashIndex), i.e. it returns a value already approximated. As well, it checks the users weights not through position.fromRewardWeight, but through rewardWeight(), which is equal to position.fromRewardWeight.mulWadDown(slashIndex), i.e. it returns a value already approximated. Indeed, Certora returns a state in which UnwindingModule.globalPoints[0x277c].totalRewardWeight=1.33e18, which is exceeded by the rewardWeight of a single user,
      UnwindingHarness.positions[0x274c].fromRewardWeight = 1.5e18. However, the precondition is satisfied, since totalRewardWeight() =
      rewardWeight(userA) = 4. Whether the state returned by Certora (i.e. where totalRewardWeight is lower than a user fromRewardWeight) is reachable, is unclear. In the report, no elaboration is provided, besides an argument about rounding error, which however does not imply the state to be reachable, but only that, if the state is reachable, then there exists a transaction that can violate the property. Our claim is that the desired property to verify is actually that point.totalRewardWeight is greater than the sum of the x.fromRewardWeight for all users.
  • Certora False / GPT-5 False: In 3 cases, both Certora and GPT-5 return False:
    • in 1 case, GPT-5 finds a counterexample similar to the state returned by Certora
    • in 2 cases, GPT-5’s counterexample is incorrect due to math errors involving scientific notation. If we instruct GPT-5 to convert numbers from scientific notation to decimal, it returns True. However, if we ask it to produce a complete trace that leads to the state provided by Certora, it returns a complete and correct trace that validates Certora’s counterexample. The states returned by Certora involve large numbers that only differs by small digits. The reason why GPT-5 was not able to find the counterexamples straight ahead is probably due to its limited math capabilities involving large numbers.